201 Assets Cash Accounts receivable, net Merchandise inventory Prepaid expenses Plant assets, net Total assets 28,837 33,e54 34,087 83,579 183,845 9,196 57,844 45,899 47,924 74,166 8,419 249,876 235,597 3,7se 289,240 474,533 489,880 S 340,980 Liabilities and Equity Accounts payable Long-term notes payable secured by $118,159 $ 67,868 44,999 mortgages on plant assets Common stock, $10 par value Retained earnings Total liabilities and equity 98,184 162,580 183,770 474 , 533 74,586 162,5e0162,5ee 58,815 489 ,880 348,980 93,148 86,37238 1. Compute the current ratio for the year ended 2017, 2016, and 2015 2. Compute the dcid-test ratio for the year ended 2017, 2016, and 2015. Complete this question by entering your answers in the tabs below. Required 1 Required 2 Compute the current ratio for the years ended Deckmber 31, 2017, 2016, and 2015. Choose Numerator Choose Denominator s Current Ratic Current ratio 2017 2016 2015 to
Answer #1
2017 2016 2015
cash 28837 33054 34087
Accounts receivable 83579 57844 45899
Mmerchandise Inventory 103045 74166 47924
Prepaid expense 9196 8419 3750
Total Current Assets 224657 173483 131660
Current Liabilities-Accounts Pay. 118159 67060 44999
Liquid Assets 112416 90898 79986
(Current Assets-Inventory-prepaid expense)
Req a:
CURRENT RATIO
NUMERATOR      / DENOMINATOR           = CURRENT RATIO
CURRENT ASSETS / CURRENT LIABILITIES   = CURRENT RATIO
2017 224657 / 118159     = 1.9
2016 173483 / 67060       = 2.59
2015 131660   / 44999       = 2.93
Req b:
ACID TEST RATIO
NUMERATOR      / DENOMINATOR           = CURRENT RATIO
LIQUID ASSETS    / CURRENT LIABILITIES   = CURRENT RATIO
2017 112416 / 118159     = 0.95
2016 90898 / 67060       = 1.36
2015 79986 / 44999       = 1.78